OpenStudy (anonymous):

The operation a *b is defined to be (a+b)(a-b) and a ◊b is defined to be (a^2) + (b^2). What is the value of (5 * 4)(5 ◊ 4).

myininaya (myininaya):

\[(5*4)=(5+4)(5-4)=(9)(1)=9\] \[(5 ◊ 4)=5^2+4^2=25+16=41\]

myininaya (myininaya):

so we have 9(41)=369

OpenStudy (anonymous):

thank you sooo much